OpenStudy (anonymous):

I WILL GIVE MEDAL Which of the following exponential functions goes through the points (1, 20) and (2, 80)? f(x) = 5(4)−x f(x) = 4(5)−x f(x) = 5(4)x f(x) = 4(5)x

OpenStudy (anonymous):

I think its the second one im not sure though

OpenStudy (error1603):

It' D

OpenStudy (anonymous):

Can you show me why you think its D?

OpenStudy (anonymous):

hes right

OpenStudy (error1603):

f(x)=4(5)x f(2)=4(52)=4(25)=100 f(x)=5(4)x f(2)=5(4)2=5(16)=80

OpenStudy (anonymous):

D was incorrect

OpenStudy (error1603):

Sorry C

OpenStudy (anonymous):

Thanks anyways im fixing to close though.

OpenStudy (error1603):

Type the wrong thing

OpenStudy (anonymous):

No I didnt because I copied and paste'd it but it came out wrong because I should have added the exponent symbol. Thanks though

OpenStudy (error1603):

No problem